find two numbers whose product is 12 such that one of the numbers is four less than eight times the other number?
Answer by algebrahouse.com(1659)

Find two numbers whose product is 12 such that one of the numbers is four less than eight times the other number?
x = one number
8x - 4 = other number
x(8x - 4) = 12 {their product is 12}
8x^2 - 4x = 12 {used distributive property}
8x^2 - 4x - 12 = 0 {subtracted 12 from both sides}
4(2x^2 - x - 3) = 0 {factored 4 out}
4(2x - 3)(x + 1) = 0 {factored into two binomials}
2x - 3 = 0 or x + 1 = 0 {set each factor equal to 0}
2x = 3 or x = -1
x = 3/2 or x = -1 {solved each equation for x}
8x - 4 = 8 or 8x - 4 = -12 {substituted into other number}

The numbers are: 3/2 and 8 or -1 and -12
www.algebrahouse.com